OpenStudy (anonymous):

Is y=(2)^x exponential growth or decay?

OpenStudy (swissgirl):

Its growing Lets plug in some values when x=1 y=2^1=2 When x=2 y=2^2=4 When x=3 y=2^3=8 When x=4 y=2^4=16 So as x gets bigger so does y so its definitely growing and increasing

OpenStudy (anonymous):

Thanks what about y=5(3)^3